The radio wave from a station broadcasting at 1.3×106 Hz has a wavelength of 230 m. What is the wave's speed? answer in m/s.
1
Expert's answer
2011-04-06T13:36:28-0400
From the formula v =λf, v = 230* 1.3×106 = 2.99×108 m/s.